# Power-based Partial Attention: Bridging Linear-Complexity and Full Attention

### v1

@@ 1 Introduction
The transformer architecture’s self-attention mechanism [23] has become the foundation of modern large language models (LLMs), enabling them to capture long-range dependencies and contextual relationships that drive remarkable performance across diverse tasks. However, this power comes at a steep cost: the standard self-attention mechanism scales quadratically with sequence length, requiring $O(L^{2})$ time complexity for a sequence of length $L$. As applications increasingly demand longer context windows—for tasks such as document understanding, code generation, and extended dialogues—this quadratic bottleneck has become critical. Doubling the sequence length quadruples the computational cost, making context windows beyond tens of thousands of tokens prohibitively expensive.
@@
A critical gap exists between $O(L)$ sliding window attention and $O(L^{2})$ full attention. Despite similar hybrid architectures, models using sliding window attention significantly underperform those with full attention, and this performance gap remains poorly understood. To address this, we introduce power-based partial attention (PPA), a parameterized attention mechanism with complexity $O(L^{1+p})$ where $0\leq p\leq 1$. This framework enables systematic exploration of the attention complexity-performance trade-off: $p=0$ corresponds to sliding window attention with linear complexity, while $p=1$ recovers full quadratic attention. By varying $p$, we can precisely characterize how model performance transitions between these extremes.
@@
  • We propose power-based partial attention (PPA), a novel attention mechanism with tunable complexity $O(L^{1+p})$ that interpolates between sliding-window (linear-complexity) attention and full (quadratic) attention.
  • We demonstrate that the performance transition follows an S-shaped curve, with rapid improvement occurring in a narrow range of $p$ values around 0.750.750.8750.875.
  • We show empirically that sub-quadratic attention ($p<1$) can achieve near-full-attention performance, challenging the necessity of $O(L^{2})$ complexity.
  • We reveal that the optimal $p$ depends on training data quantity, with more extensive training potentially lowering the required complexity.
@@ 2 Methodology
We can view this as the density of attention, such that the average amount of attention per token is then $\frac{1}{\sqrt{i}}$ for token $i$. Then, we can integrate the density to calculate the total number of attended tokens as $\int_{0}^{i}\frac{1}{\sqrt{j}}dj=2j^{\frac{1}{2}}\big|_{0}^{i}=2i^{\frac{1}{2}}\approx O(\sqrt{i})$. Performing another integration to sum over all tokens yields $\int_{0}^{L}i^{\frac{1}{2}}di=\frac{2}{3}L^{\frac{3}{2}}\approx O(L^{\frac{3}{2}})$ for the whole sequence.
Using this strategy, the attended relative offsets (backward from the current token) can be taken from the squared values $1,4,9,16,25,36,\ldots$ where $1$ means one position back, $4$ means four positions back, and so on. This pattern preserves the $\sqrt{L}$ scaling for each token. For example, for a sequence of length $25$, the $25$th token attends to only $\sqrt{25}=5$ tokens. We can extend this methodology to general $p=1/n$, where $n$ is a positive integer. For instance, $p=\frac{1}{3}$ corresponds to attending to positions $1,8,27,64,125,216,\ldots$ (cubic positions). We can generalize this methodology further for any $0<p<1$, such that whether position $j$ is an attended position can be determined by checking if $\lfloor j^{p}\rfloor$ increments by 1, where $\lfloor\cdot\rfloor$ denotes the floor function.
@@
In this study, we define the PPA causal mask as the union of sliding window attention and the power-based offsets described above. Consider token positions $\{1,\ldots,L\}$, a sliding window size $w\geq 1$, and a power parameter $p\in[0,1]$. For $p\in(0,1]$, we first define the set of attended relative offsets
$\mathcal{S}_{p}\;=\;\Big\{j\in\mathbb{N}:\lfloor j^{p}\rfloor-\lfloor(j-1)^{p}\rfloor=1\Big\}.$
Then, for each query position $i$, the allowed keys are
$\mathcal{K}_{p}(i)\;=\;\Big\{k<i:(i-k)\leq w\ \text{or}\ \ (i-k)\in\mathcal{S}_{p}\Big\},$
and the causal mask is $M_{p}(i,k)=1$ if $k\in\mathcal{K}_{p}(i)$ and $M_{p}(i,k)=0$ otherwise. When $p=1$, this reduces to full causal attention, and when $p=0$ we recover sliding window attention.
This yields $O(L^{1+p})$ attention overall: token $i$ attends to $O(i^{p})$ power-based links (plus $O(w)$ local links), so summing over $i=1,\ldots,L$ gives $\sum_{i}O(i^{p})=O(L^{1+p})$ when $w$ is fixed.
@@ 3 Implementation and Experimentation
Using the power-based attention mechanism described above, we can vary $p$ to observe the transition from linear scaling ($O(L)$) to full attention ($O(L^{2})$). To use this mechanism more effectively, we combine it with sliding window attention. Since sliding window attention has $O(L)$ complexity, it does not affect the overall asymptotic scaling, but it is crucial for generating coherent text by making recent tokens visible to the model. By taking the union of the power-based attention mechanism and sliding window attention, and varying $p$ from 0 to $1$, we can directly observe how the model transitions from pure sliding window attention to full attention.
@@ 4 Results and Discussion
Conversely, the widely used full attention is not necessarily optimal either, as we observe diminishing returns at large $p$. However, it is important to keep in mind that although $O(L^{1+p})$ theoretically scales slower than $O(L^{2})$, in practice, the highly optimized full attention implementation using FlashAttention is likely more efficient than sub-quadratic variants with less-optimized kernels [8].

### v2

@@ 1 Introduction
The transformer architecture’s self-attention mechanism has become the foundation of modern large language models (LLMs), enabling them to capture long-range dependencies and contextual relationships that drive remarkable performance across diverse tasks. However, this power comes at a steep cost: the standard self-attention mechanism scales quadratically with sequence length, requiring $O(L^{2})$ time and memory for a sequence of length $L$. As applications increasingly demand longer context windows—for tasks such as document understanding, code generation, and extended dialogues—this quadratic bottleneck has become critical. Doubling the sequence length quadruples both computational cost and memory requirements, making context windows beyond tens of thousands of tokens prohibitively expensive.
@@
A critical gap exists between $O(L)$ sliding window attention and $O(L^{2})$ full attention. Despite similar hybrid architectures, models using sliding window attention significantly underperform those with full attention, and this performance gap remains poorly understood. To address this, we introduce power-based partial attention (PPA), a parameterized attention mechanism with complexity $O(L^{1+p})$ where $0\leq p\leq 1$. This framework enables systematic exploration of the attention complexity-performance trade-off: $p=0$ corresponds to sliding window attention with linear complexity, while $p=1$ recovers full quadratic attention. By varying $p$, we can precisely characterize how model performance transitions between these extremes. Recently and concurrent with our work, PowerAttention [5] explored a related power-based sub-quadratic attention direction. While both efforts share the high-level motivation, our focus here is a simple causal masking construction (incremental stride selection combined with sliding window attention) and a controlled sweep over $p$ to directly quantify the resulting complexity–performance transition.
@@
  • We introduce a simple, parameterized causal attention-mask family with tunable complexity $O(L^{1+p})$ that continuously interpolates between sliding-window (linear-complexity) attention and full (quadratic) attention.
  • We provide a controlled empirical characterization of the complexity–performance relationship as a function of $p$, revealing an S-shaped transition with rapid improvement in a narrow range (around $p\approx 0.75$$0.875$ in our experiments).
  • We show that sub-quadratic attention ($p<1$) can achieve near-full-attention performance on MATH500 and GSM8k after task-specific adaptation.
  • We find that the optimal $p$ depends on training data quantity and prompt structure, suggesting that additional training can lower the required attention complexity.
@@ 2 Methodology
We can view this as the density of attention, such that the average amount of attention per token is then $\frac{1}{\sqrt{i}}$ for token $i$. Then, we can integrate the density to calculate the total number of attended tokens as $\int_{0}^{i}\frac{1}{\sqrt{j}}dj=\frac{1}{2}j^{\frac{1}{2}}\big|_{0}^{i}=\frac{1}{2}i^{\frac{1}{2}}\approx O(\sqrt{i})$. Performing another integration to sum over all tokens yields $\int_{0}^{L}i^{\frac{1}{2}}di=\frac{2}{3}L^{\frac{3}{2}}\approx O(L^{\frac{3}{2}})$ for the whole sequence.
Using this strategy, the token positions to attend to can be taken from the squared values $1,4,9,16,25,36,\ldots$ where $1$ is the current token position, $4$ is three positions back, and so on. This pattern preserves the $\sqrt{L}$ scaling for each token. For example, for a sequence of length $25$, the $25$th token attends to only $\sqrt{25}=5$ tokens. We can extend this methodology to general $p=1/n$, where $n$ is a positive integer. For instance, $p=\frac{1}{3}$ corresponds to attending to positions $1,8,27,64,125,216,\ldots$ (cubic positions). We can generalize this methodology for any $0<p<1$, such that whether position $j$ is an attended position can be determined by checking if $\lfloor j^{p}\rfloor$ increments by 1, where $\lfloor\cdot\rfloor$ denotes the floor function.
@@ 3 Implementation and Experimentation
Using the power-based attention mechanism described above, we can vary $p$ to observe the transition from sliding window attention ($O(L)$) to full attention ($O(L^{2})$). To use this mechanism more effectively, we combine it with sliding window attention. Since sliding window attention has $O(L)$ complexity, it does not affect the overall asymptotic scaling, but it is crucial for generating coherent text by making recent tokens visible to the model. By taking the union of the power-based attention mechanism and sliding window attention, and varying $p$ from 0 to $1$, we can directly observe how the model transitions from pure sliding window attention to full attention.
@@ 4 Results and Discussion
Conversely, the widely used full attention is not necessarily optimal either, as we observe diminishing returns at large $p$. However, it is important to keep in mind that although $O(L^{1+p})$ theoretically scales slower than $O(L^{2})$, in practice, the highly optimized full attention implementation using FlashAttention is likely more efficient than sub-quadratic variants with less-optimized kernels.